Diffusion Morning Report 05/07/04

ted berg tberg at snf.stanford.edu
Fri May 7 06:58:37 PDT 2004

Good Morning All,
    Looking good this morning.  Nitride is back together and coated 
waiting for a test.  LTO is being vacuumed off and scheduled for
Pull and clean on Monday.  TEOS 2 is now fully assembled including 
trap-testing. Power supply for Tylan 1 is on order. Burn box cleaned this
morning. Everything else looks OK. 

Diffusion 	Date: 	5/7/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	down 	I/O error-bad PS on Order 	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	MFC is installed testing-trap installed 	 
Tylan 8-SiGe 	Up 	  	6.6/30
Tylan 9-Poly 	Up 	  	27.3/30
Tylan 10-Nitride 	down 	coat done- needs test 	0/30
Tylan 11-BPSG-LTO 	Up 	  	30.13/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	11.2/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 550 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040507/d0012d09/attachment.html>

More information about the diffusion mailing list